OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

OpenStudy (blurbendy):

use the formula y - y1 = slope*(x - x1) Substitute for y1 and x1 y - 4 = -1(x + 1) y - 4 = -x - 1 y = -x - 3